OpenStudy (anonymous):

A salesperson earns a commission based on the number and type of vehicle sold. A person selling 6 cars and 3 trucks earns $4,800. A person selling 8 cars and 1 truck earns $4,600. How much does a salesperson earn for selling 2 cars and 3 trucks? a. $2,500 b. $2,700 c. $2,800 d. $3,000

OpenStudy (anonymous):

@texaschic101 I think its D. am I correct ?

OpenStudy (texaschic101):

6c + 3t = 4800 8c + t = 4600 --->(-3) ------------- 6c + 3t = 4800 -24c - 3t = - 13,800 (result of multiplying by -3) --------------add - 18c = - 9000 c = 500 6c + 3t = 4800 6(500) + 3t = 4800 3000 + 3t = 4800 3t = 4800 - 3000 3t = 1800 t = 600 6c + 3t = 4800 6(500) + 3(600) = 4800 3000 + 1800 = 4800 4800 = 4800 (correct) 2c + 3t = 2(500) + 3(600) = 1000 + 1800 = 2800 <=== the answer I am getting is C

OpenStudy (andrewthecookie):

Its C nice despriction texas

OpenStudy (texaschic101):

thank you :)

OpenStudy (anonymous):

Thank you ! I see where I messed up :)